OpenStudy (hba):

one root of the equation x^2+px+q=0 is 2-3i.The value of p and q is :

OpenStudy (hba):

@ash2326

OpenStudy (sirm3d):

the other root is the conjugate of 2-3i. the sum of the roots is -b/a while the product of the roots is c/a. can you get the sum and product of the two complex roots?

OpenStudy (hba):

I know what is S.O.R And P.O.R but i do not know how to get them ?

OpenStudy (sirm3d):

one root is r1 = 2 - 3i, the other root is r2 = 2 + 3i. if you get the sum of the roots r1 + r2, you get 4

OpenStudy (hba):

Wait

OpenStudy (sirm3d):

-b/a = 4 with a = 1 (coefficient of x^2), we get b = -4. can you do the product of roots part?

OpenStudy (hba):

How did you get a=1

OpenStudy (sirm3d):

x^2 = 1 * x^2, so there, the coefficient is 1